OpenStudy (babybritt002):

you have 1500.g of bleach solution. The percent by mass of the solute sodium hypochlorite (NaOCI) is 3.62%. How many grams of NaOCI are in the solution? Please show work

OpenStudy (babybritt002):

@MTALHAHASSAN2

OpenStudy (mtalhahassan2):

ok let me see

OpenStudy (babybritt002):

@MTALHAHASSAN2 ok

OpenStudy (mtalhahassan2):

3.61 = mass NaClO x 100/ 1500 mass NaClO = 54.2 g 2.65 = 50.0 x 100 / mass solution mass solution = 1887 g % v/v= 35 x 100 / 190 = 18.4 moles glucose = 5.10 g / 180 g/mol=0.0283 M = 0.0283 / 0.1005 L=0.282

OpenStudy (mtalhahassan2):

here you go

OpenStudy (mtalhahassan2):

i hope it will help

OpenStudy (mtalhahassan2):

@babybritt002

OpenStudy (babybritt002):

@MTALHAHASSAN2 how did you get 3.61 where its 3.62?

OpenStudy (mtalhahassan2):

oh Is it 3.62?? my bad

OpenStudy (mtalhahassan2):

i apologize for that

OpenStudy (babybritt002):

@MTALHAHASSAN2 yea so can you do it for me ?

OpenStudy (babybritt002):

Or do you not know how to ? @MTALHAHASSAN2